# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 3.3 Existence and uniqueness of weak solution
$\qquad\qquad\qquad+\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\gamma\Big{(}C_{1}\,(\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{1}{c_{a}}\,\gamma\,r\Big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}$
$\quad\leq\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\bigg{\{}\kappa_{2}\,\widetilde{C}(1+C_{1}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\boldsymbol{f}\|_{0,\Omega})+\dfrac{\gamma^{2}}{c_{a}}(1+\kappa_{2}\,\widetilde{C}\,r)\bigg{\}}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega},$
which together with (3.15) and the Banach fixed-point theorem yields that $\mathbf{T}$ has a unique fixed point in $\mathrm{W}$. Finally, (3.16) is derived analogously to the estimates in (3.7) and (3.9), which completes the proof.
@@
The operator $\mathbf{T}$ (see (3.5)) could be also defined, for example $\mathbf{T}:\mathbf{W}\to\mathbf{W}$, with $\mathbf{W}:=\Big{\{}(\widehat{\boldsymbol{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{L}^{2}(\Omega)\,:\quad\|\widehat{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\widehat{p}\|_{0,\Omega}\leq r\Big{\}}$ and $\mathbf{T}(\widehat{\boldsymbol{\sigma}},\widehat{p}):=(R_{1}(\widehat{p}),S_{2}(\widehat{\boldsymbol{\sigma}},\widehat{p}))=(\boldsymbol{\sigma},p)$, with $R_{1}$ and $S_{2}$ defined as in (3.3) and (3.4), respectively.
@@ 4.1 Finite element spaces and Galerkin scheme
Let us consider a regular partition $\mathcal{T}_{h}$ of $\bar{\Omega}$ made up of triangles $K$ (in $\mathbb{R}^{2}$) or tetrahedra $K$ (in $\mathbb{R}^{3}$) of diameter $h_{K}$, and denote the mesh size by $h:=\mathop{\rm max}\limits\{h_{K}:\ K\in\mathcal{T}_{h}\}$. Given an integer $\ell\geq 0$ and $K\in\mathcal{T}_{h}$, we first let $\mathrm{P}_{\ell}(K)$ be the space of polynomials of degree $\leq\ell$ defined on $K$, whose vector and tensor versions are denoted $\mathbf{P}_{\ell}(K)\,:=\,[\mathrm{P}_{\ell}(K)]^{d}$ and $\mathbb{P}_{\ell}(K)\,=\,[\mathrm{P}_{\ell}(K)]^{d\times d}$, respectively. Also, we let $\mathbf{RT}_{\ell}(K)\,:=\,\mathbf{P}_{\ell}(K)\oplus\mathrm{P}_{\ell}(K)\,\boldsymbol{x}$ be the local Raviart–Thomas space of order $\ell$ defined on $K$, where $\boldsymbol{x}$ stands for a generic vector in $\mathbb{R}^{d}$, and denote by $\mathbb{RT}_{k}(K)$ the tensor-valued counterpart of this space.
@@
$\mathbf{B}_{k}(K):=\begin{cases}\mathbf{curl}^{\mathrm{t}}(b_{K}\mathrm{P}_{k}(K))&\textrm{in}\quad\mathbb{R}^{2},\\ \nabla\times(b_{K}\mathbf{P}_{k}(K))&\textrm{in}\quad\mathbb{R}^{3},\end{cases}$
@@
We recall the classical PEERSk elements (cf. [5]) to define the discrete subspaces for the stress tensor $\boldsymbol{\sigma}$, the displacement $\boldsymbol{u}$, and the rotation tensor $\boldsymbol{\rho}$
$\mathbb{H}^{\boldsymbol{\sigma}}_{h}$ :=$:=\left\{{\boldsymbol{\tau}}_{h}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega):\quad{\boldsymbol{\tau}}_{h}|_{K}\in\mathbb{RT}_{k}(K)\oplus[\mathbf{B}_{k}(K)]^{d}\quad\forall\,K\in\mathcal{T}_{h}\right\},$
(4.1) $\mathbf{H}^{\boldsymbol{u}}_{h}$ :=$:=\left\{\boldsymbol{v}_{h}\in\mathbf{L}^{2}(\Omega):\quad\boldsymbol{v}_{h}|_{K}\in\textbf{P}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
$\mathbb{H}^{\boldsymbol{\rho}}_{h}$ :=$:=\left\{{\boldsymbol{\eta}}_{h}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\cap\mathbb{C}(\overline{\Omega})\quad\mathrm{and}\quad{\boldsymbol{\eta}}_{h}|_{K}\in\mathbb{P}_{k+1}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
and the following estimates are proven for the PEERSk elements (cf. [41, Remark 3.3])
(4.2a) $\sup_{\boldsymbol{0}\neq\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}}\frac{b(\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))}{\|\boldsymbol{\tau}_{h}\|_{\mathbf{div};\Omega}}$ $\geq\beta^{*}(\|\boldsymbol{v}_{h}\|_{0,\Omega}+\|{\boldsymbol{\eta}}_{h}\|_{0,\Omega})\quad\forall\,(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})\in\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h},$
(4.2b) $\displaystyle a(\boldsymbol{\tau}_{h},\boldsymbol{\tau}_{h})$ $\geq c_{a}\|\boldsymbol{\tau}_{h}\|_{\mathbf{div};\Omega}^{2}\quad\forall\,\boldsymbol{\tau}_{h}\in\mathbb{V}_{h},$
@@
(4.3) $\mathbb{V}_{h}:=\displaystyle\Big{\{}\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}:\quad b(\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))=0\quad\forall\,(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})\in\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\Big{\}}.$
Additionally, for ${\boldsymbol{\varphi}}$ and the pressure $p$, we define the FE subspaces
$\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}$ :=$:=\left\{{\boldsymbol{\psi}}_{h}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega):\ {\boldsymbol{\psi}}_{h}|_{K}\in\mathbf{RT}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
(4.4) $\mathrm{H}^{p}_{h}$ :=$:=\left\{q_{h}\in\mathrm{L}^{2}(\Omega):\ q_{h}|_{K}\in\mathrm{P}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
@@
(4.5) $\sup_{\boldsymbol{0}\neq{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}}\frac{\widetilde{b}({\boldsymbol{\psi}}_{h},q_{h})}{\|{\boldsymbol{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}}\geq\widetilde{\beta}^{*}\|q_{h}\|_{0,\Omega}\quad\forall\,q_{h}\in\mathrm{H}^{p}_{h}.$
Note that it is of course possible to consider other conforming and inf-sup stable spaces such as Arnold–Falk–Winther and Brezzi–Douglas–Marini instead of (4.1) and (4.1), respectively. The Galerkin scheme for (2.2) reads: find $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, such that:
(4.6) $\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})&+\quad\displaystyle b(\boldsymbol{\tau}_{h},(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))&+\quad c(\boldsymbol{\tau}_{h},p_{h})&=\displaystyle H(\boldsymbol{\tau}_{h})&\forall\,\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma}_{h},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))&&&=\displaystyle F(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})&\forall(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})\in\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h},\\[4.30554pt] \displaystyle\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}({\boldsymbol{\varphi}}_{h},{\boldsymbol{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}}_{h},p_{h})&&=\displaystyle\widetilde{H}({\boldsymbol{\psi}}_{h})&\forall\,{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&-\quad c(\boldsymbol{\sigma}_{h},q_{h})&=\displaystyle G(q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$
@@ 4.2 Analysis of the discrete problem
Next, for a fixed $\widehat{p}_{h}$ in $\mathrm{W}_{h}$, we have that the bilinear form $\widetilde{a}_{\boldsymbol{\sigma}_{h},p_{h}}$ satisfies
(4.7) $\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h}}({\boldsymbol{\psi}}_{h},{\boldsymbol{\psi}}_{h})\geq\kappa_{1}\|{\boldsymbol{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}^{2}\quad\forall\,{\boldsymbol{\psi}}_{h}\in\mathbf{V}_{h},$
@@
(4.8) $\begin{array}[]{ll}\displaystyle\mathbf{V}_{h}:=\displaystyle\Big{\{}{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}:\quad\widetilde{b}({\boldsymbol{\psi}}_{h},q_{h})=0\quad\forall\,q_{h}\in\mathrm{H}^{p}_{h}\Big{\}}.\end{array}$
Additionally, we define the discrete operators $\mathbf{R}_{h}:\mathrm{W}_{h}\subseteq\mathrm{H}^{p}_{h}\to\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$ and $\mathbf{S}_{h}:\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathrm{W}_{h}\to\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, respectively, by
$\mathbf{R}_{h}(\widehat{p}_{h}):=\big{(}R_{1,h}(\widehat{p}_{h}),(R_{2,h}(\widehat{p}_{h}),R_{3,h}(\widehat{p}_{h}))\big{)}=(\boldsymbol{\sigma}_{h},(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))\quad\forall\,\widehat{p}_{h}\in\mathrm{W}_{h},$
where $(\boldsymbol{\sigma}_{h},(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$ is the unique solution of
(4.9) $\begin{array}[]{rllll}\displaystyle a(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})&+\quad\displaystyle b(\boldsymbol{\tau}_{h},(\boldsymbol{u}_{h},\boldsymbol{\rho}_{h}))&=\displaystyle H(\boldsymbol{\tau}_{h})-c(\boldsymbol{\tau}_{h},\widehat{p}_{h})&\forall\,\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\boldsymbol{\sigma}_{h},(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h}))&&=\displaystyle F(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})&\forall\,(\boldsymbol{v}_{h},{\boldsymbol{\eta}}_{h})\in\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h},\end{array}$
@@
$\mathbf{S}_{h}(\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h}):=\big{(}S_{1,h}(\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h}),S_{2,h}(\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h})\big{)}=({\boldsymbol{\varphi}}_{h},p_{h})\quad\forall\,(\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathrm{W}_{h},$
where $({\boldsymbol{\varphi}}_{h},p_{h})$ is the unique tuple in $\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\times\mathrm{H}^{p}_{h}$ such that
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\boldsymbol{\sigma}}_{h},\widehat{p}_{h}}({\boldsymbol{\varphi}}_{h},{\boldsymbol{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}}_{h},p_{h})&=\displaystyle\widetilde{H}({\boldsymbol{\psi}}_{h})&\forall\,{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\boldsymbol{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&=\displaystyle G(q_{h})+c(\widehat{\boldsymbol{\sigma}}_{h},q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$
Employing properties (4.2a), (4.2b), (4.5), (4.7) and (2.9) and proceeding exactly as for the continuous case (Lemmas 3.1 and 3.3), it can be easily deduced that both operators are well-defined. Then, analogously to the continuous case, we define the following fixed-point operator
(4.10) $\mathbf{T}_{h}:\mathrm{W}_{h}\subseteq\mathrm{H}^{p}_{h}\to\mathrm{H}^{p}_{h},\quad\widehat{p}_{h}\mapsto\mathbf{T}_{h}(\widehat{p}_{h}):=S_{2,h}(R_{1,h}(\widehat{p}_{h}),\widehat{p}_{h}),$

### v2

@@ 3.3 Existence and uniqueness of weak solution
$\qquad+\dfrac{2\kappa_{2}\,\widetilde{C}}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\gamma\Big{(}C_{1}\,(\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})+\dfrac{1}{c_{a}}\,\gamma\,r\Big{)}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega}$
$\leq\dfrac{2}{\min\{\widetilde{\gamma},\,\kappa_{1}\}}\,\bigg{\{}\kappa_{2}\,\widetilde{C}(1+C_{1}\gamma)(\|g\|_{0,\Omega}+\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}+\|\bm{f}\|_{0,\Omega})$
$\qquad\qquad\qquad\qquad+\dfrac{\gamma^{2}}{c_{a}}(1+\kappa_{2}\,\widetilde{C}\,r)\bigg{\}}\|\widehat{p}_{1}-\widehat{p}_{2}\|_{0,\Omega},$
which together with (3.15) and the Banach fixed-point theorem yields that $\mathbf{T}$ has a unique fixed point in $\mathrm{W}$. Finally, (3.8) is derived analogously to the estimates in (3.7) and (3.9), which completes the proof. ∎
@@
The operator $\mathbf{T}$ (see (3.5)) could be also defined, for example $\mathbf{T}:\mathbf{W}\to\mathbf{W}$, with $\mathbf{W}:=\Big{\{}(\widehat{\bm{\sigma}},\widehat{p})\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{L}^{2}(\Omega)\,:\quad\|\widehat{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\widehat{p}\|_{0,\Omega}\leq r\Big{\}}$ and $\mathbf{T}(\widehat{\bm{\sigma}},\widehat{p}):=(R_{1}(\widehat{p}),S_{2}(\widehat{\bm{\sigma}},\widehat{p}))=(\bm{\sigma},p)$, with $R_{1}$ and $S_{2}$ defined as in (3.3) and (3.4), respectively.
@@ 4.1 Finite element spaces and Galerkin scheme
Let us consider a regular partition $\mathcal{T}_{h}$ of $\bar{\Omega}$ made up of triangles $K$ (in $\mathbb{R}^{2}$) or tetrahedra $K$ (in $\mathbb{R}^{3}$) of diameter $h_{K}$, and denote the mesh size by $h:=\mathop{\rm max}\limits\{h_{K}:\ K\in\mathcal{T}_{h}\}$. Given an integer $\ell\geq 0$ and $K\in\mathcal{T}_{h}$, we first let $\mathrm{P}_{\ell}(K)$ be the space of polynomials of degree $\leq\ell$ defined on $K$, whose vector and tensor versions are denoted $\mathbf{P}_{\ell}(K)\,:=\,[\mathrm{P}_{\ell}(K)]^{d}$ and $\mathbb{P}_{\ell}(K)\,=\,[\mathrm{P}_{\ell}(K)]^{d\times d}$, respectively. Also, we let $\mathbf{RT}_{\ell}(K)\,:=\,\mathbf{P}_{\ell}(K)\oplus\mathrm{P}_{\ell}(K)\,\bm{x}$ be the local Raviart–Thomas space of order $\ell$ defined on $K$, where $\bm{x}$ stands for a generic vector in $\mathbb{R}^{d}$, and denote by $\mathbb{RT}_{k}(K)$ the tensor-valued counterpart of this space.
@@
$\mathbf{B}_{k}(K):=\begin{cases}{\mathrm{curl}(b_{K})\mathrm{P}_{k}(K)}&\textrm{in}\quad\mathbb{R}^{2},\\ \nabla\times(b_{K}\mathbf{P}_{k}(K))&\textrm{in}\quad\mathbb{R}^{3},\end{cases}$
@@
We recall the classical PEERSk elements (cf. [5]) to define the discrete subspaces for the stress tensor $\bm{\sigma}$, the displacement ${\bm{u}}$, and the rotation tensor $\bm{\rho}$
$\mathbb{H}^{\bm{\sigma}}_{h}$ :=$:=\left\{{\bm{\tau}}_{h}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega):\quad{\bm{\tau}}_{h}|_{K}\in\mathbb{RT}_{k}(K)\oplus[\mathbf{B}_{k}(K)]^{d}\quad\forall\,K\in\mathcal{T}_{h}\right\},$
$\mathbf{H}^{{\bm{u}}}_{h}$ :=$:=\left\{\mathbf{v}_{h}\in\mathbf{L}^{2}(\Omega):\quad\mathbf{v}_{h}|_{K}\in\textbf{P}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$ (4.1)
$\mathbb{H}^{\bm{\rho}}_{h}$ :=$:=\left\{{\bm{\eta}}_{h}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\cap\mathbb{C}(\overline{\Omega})\quad\mathrm{and}\quad{\bm{\eta}}_{h}|_{K}\in\mathbb{P}_{k+1}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
where $\mathbb{C}(\overline{\Omega})$ denotes the space of continuous tensor fields, and the following estimates are proven for the PEERSk elements (cf. [42, Remark 3.3])
$\sup_{\bm{0}\neq\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}}\frac{b(\bm{\tau}_{h},(\mathbf{v}_{h},{\bm{\eta}}_{h}))}{\|\bm{\tau}_{h}\|_{\mathbf{div};\Omega}}$ $\geq\beta^{*}(\|\mathbf{v}_{h}\|_{0,\Omega}+\|{\bm{\eta}}_{h}\|_{0,\Omega})\quad\forall\,(\mathbf{v}_{h},{\bm{\eta}}_{h})\in\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h},$ (4.2a)
$\displaystyle a(\bm{\tau}_{h},\bm{\tau}_{h})$ $\geq c_{a}\|\bm{\tau}_{h}\|_{\mathbf{div};\Omega}^{2}\quad\forall\,\bm{\tau}_{h}\in\mathbb{V}_{h},$ (4.2b)
@@
$\mathbb{V}_{h}:=\displaystyle\Big{\{}\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}:\quad b(\bm{\tau}_{h},(\mathbf{v}_{h},{\bm{\eta}}_{h}))=0\quad\forall\,(\mathbf{v}_{h},{\bm{\eta}}_{h})\in\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\Big{\}}.$
Additionally, for ${\bm{\varphi}}$ and the pressure $p$, we define the FE subspaces
$\mathbf{H}^{{\bm{\varphi}}}_{h}$ :=$:=\left\{{\bm{\psi}}_{h}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega):\ {\bm{\psi}}_{h}|_{K}\in\mathbf{RT}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$
$\mathrm{H}^{p}_{h}$ :=$:=\left\{q_{h}\in\mathrm{L}^{2}(\Omega):\ q_{h}|_{K}\in\mathrm{P}_{k}(K)\quad\forall\,K\in\mathcal{T}_{h}\right\},$ (4.3)
@@
$\sup_{\bm{0}\neq{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h}}\frac{\widetilde{b}({\bm{\psi}}_{h},q_{h})}{\|{\bm{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}}\geq\widetilde{\beta}^{*}\|q_{h}\|_{0,\Omega}\quad\forall\,q_{h}\in\mathrm{H}^{p}_{h}.$ (4.4)
Note that it is of course possible to consider other conforming and inf-sup stable spaces such as Arnold–Falk–Winther and Brezzi–Douglas–Marini instead of (4.1) and (4.1), respectively. The Galerkin scheme for (2.2) reads: find $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, such that:
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma}_{h},\bm{\tau}_{h})&+\quad\displaystyle b(\bm{\tau}_{h},({\bm{u}}_{h},\bm{\rho}_{h}))&+\quad c(\bm{\tau}_{h},p_{h})&=\displaystyle H(\bm{\tau}_{h})&\forall\,\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\bm{\sigma}_{h},(\mathbf{v}_{h},{\bm{\eta}}_{h}))&&&=\displaystyle F(\mathbf{v}_{h},{\bm{\eta}}_{h})&\forall(\mathbf{v}_{h},{\bm{\eta}}_{h})\in\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h},\\[4.30554pt] \displaystyle\widetilde{a}_{\bm{\sigma}_{h},p_{h}}({\bm{\varphi}}_{h},{\bm{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\bm{\psi}}_{h},p_{h})&&=\displaystyle\widetilde{H}({\bm{\psi}}_{h})&\forall\,{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&-\quad c(\bm{\sigma}_{h},q_{h})&=\displaystyle{\tilde{F}}(q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$ (4.5)
@@ 4.2 Analysis of the discrete problem
Next, for a fixed $\widehat{p}_{h}$ in $\mathrm{W}_{h}$, we have that the bilinear form $\widetilde{a}_{\bm{\sigma}_{h},p_{h}}$ satisfies
$\widetilde{a}_{\widehat{\bm{\sigma}}_{h},\widehat{p}_{h}}({\bm{\psi}}_{h},{\bm{\psi}}_{h})\geq\kappa_{1}\|{\bm{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}^{2}\quad\forall\,{\bm{\psi}}_{h}\in\mathbf{V}_{h},$ (4.6)
@@
$\mathbf{V}_{h}:=\displaystyle\Big{\{}{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h}:\quad\widetilde{b}({\bm{\psi}}_{h},q_{h})=0\quad\forall\,q_{h}\in\mathrm{H}^{p}_{h}\Big{\}}.$
Additionally, we define the discrete operators $\mathbf{R}_{h}:\mathrm{W}_{h}\subseteq\mathrm{H}^{p}_{h}\to\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$ and $\mathbf{S}_{h}:\mathbb{H}^{\bm{\sigma}}_{h}\times\mathrm{W}_{h}\to\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$, respectively, by
$\mathbf{R}_{h}(\widehat{p}_{h}):=\big{(}R_{1,h}(\widehat{p}_{h}),(R_{2,h}(\widehat{p}_{h}),R_{3,h}(\widehat{p}_{h}))\big{)}=(\bm{\sigma}_{h},({\bm{u}}_{h},\bm{\rho}_{h}))\quad\forall\,\widehat{p}_{h}\in\mathrm{W}_{h},$
where $(\bm{\sigma}_{h},({\bm{u}}_{h},\bm{\rho}_{h}))\in\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$ is the unique solution of
$\begin{array}[]{rllll}\displaystyle a(\bm{\sigma}_{h},\bm{\tau}_{h})&+\quad\displaystyle b(\bm{\tau}_{h},({\bm{u}}_{h},\bm{\rho}_{h}))&=\displaystyle H(\bm{\tau}_{h})-c(\bm{\tau}_{h},\widehat{p}_{h})&\forall\,\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h},\\[4.30554pt] \displaystyle b(\bm{\sigma}_{h},(\mathbf{v}_{h},{\bm{\eta}}_{h}))&&=\displaystyle F(\mathbf{v}_{h},{\bm{\eta}}_{h})&\forall\,(\mathbf{v}_{h},{\bm{\eta}}_{h})\in\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h},\end{array}$ (4.7)
@@
$\mathbf{S}_{h}(\widehat{\bm{\sigma}}_{h},\widehat{p}_{h}):=\big{(}S_{1,h}(\widehat{\bm{\sigma}}_{h},\widehat{p}_{h}),S_{2,h}(\widehat{\bm{\sigma}}_{h},\widehat{p}_{h})\big{)}=({\bm{\varphi}}_{h},p_{h})\quad\forall\,(\widehat{\bm{\sigma}}_{h},\widehat{p}_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathrm{W}_{h},$
where $({\bm{\varphi}}_{h},p_{h})$ is the unique tuple in $\mathbf{H}^{{\bm{\varphi}}}_{h}\times\times\mathrm{H}^{p}_{h}$ such that
$\begin{array}[]{rllll}\displaystyle\widetilde{a}_{\widehat{\bm{\sigma}}_{h},\widehat{p}_{h}}({\bm{\varphi}}_{h},{\bm{\psi}}_{h})&+\quad\displaystyle\widetilde{b}({\bm{\psi}}_{h},p_{h})&=\displaystyle\widetilde{H}({\bm{\psi}}_{h})&\forall\,{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h},\\[4.30554pt] \displaystyle\widetilde{b}({\bm{\varphi}}_{h},q_{h})&-\quad\widetilde{c}(p_{h},q_{h})&=\displaystyle{\tilde{F}}(q_{h})+c(\widehat{\bm{\sigma}}_{h},q_{h})&\forall\,q_{h}\in\mathrm{H}^{p}_{h}.\end{array}$
Employing properties (4.2a), (4.2b), (4.4), (4.6) and (2.9) and proceeding exactly as for the continuous case (Lemmas 3.1 and 3.2), it can be easily deduced that both operators are well-defined. Then, analogously to the continuous case, we define the following fixed-point operator
$\mathbf{T}_{h}:\mathrm{W}_{h}\subseteq\mathrm{H}^{p}_{h}\to\mathrm{H}^{p}_{h},\quad\widehat{p}_{h}\mapsto\mathbf{T}_{h}(\widehat{p}_{h}):=S_{2,h}(R_{1,h}(\widehat{p}_{h}),\widehat{p}_{h}),$ (4.8)